OpenStudy (anonymous):

Four students are working in a group and are asked to find the length of BC. Which two students set up the problem correctly?

OpenStudy (anonymous):

\[Moe: x = \sqrt{(5 - 1)^{2}+ (1 - 4)^{2}}\] Jimmy: x2 = 32 + 42 Alice: x = 3 + 4 Jane: x = \[Jane: x =\sqrt{(5-1^{2} - (1-4)^{2}} \]

OpenStudy (anonymous):

@jojor_124 what grade is this

OpenStudy (anonymous):

As in school? 10th grade. Geometry, and stuff.

OpenStudy (anonymous):

Hmmm

OpenStudy (anonymous):

Look I really want to help but I'm only in 8th grade algebra right now, this upcoming year I'm going into 9th grade geometry, I'm not really familiar with this stuff :/ I could try to find someone smarter to help you though lol

OpenStudy (anonymous):

Kk c:

OpenStudy (anonymous):

@DangerousJesse

OpenStudy (anonymous):

Do you understand this @DangerousJesse bc I really don't

OpenStudy (dangerousjesse):

For the most part, yes :) Let me work it out right quick

OpenStudy (dangerousjesse):

Have you worked through the equations already?

OpenStudy (anonymous):

I don't know where @jojor_124 went

OpenStudy (dangerousjesse):

Well, Moe says that \(x=5\)

OpenStudy (dangerousjesse):

Jimmy says \(x=37\)

OpenStudy (dangerousjesse):

Alice says \(x=7\)

OpenStudy (dangerousjesse):

And Jane says \(x= \sqrt{7}\)

OpenStudy (anonymous):

Wow you're good @DangerousJesse

OpenStudy (anonymous):

Oh my._. So which two did it correctly?

OpenStudy (dangerousjesse):

Well, the distance between (4,1) and (1,5) is 5, so you tell me :P

OpenStudy (anonymous):

So would it be Moe and Jimmy, or Moe and Alice?

OpenStudy (dangerousjesse):

I only got Moe..

OpenStudy (anonymous):

I'm gonna go with Moe andAlice. Let's pray I got dis.

OpenStudy (dangerousjesse):

Good luck >-<